Scott is 22 years older than Sun Jung, In 7 years, Scott will be twice as old as Sun Jung. Give the current age of Scott and Sun Jung.
Answer by ankor@dixie-net.com(22740)

Write an equation for each statement:
:
Scott is 22 years older than Sun Jung,
S = J + 22
:
In 7 years, we have:
(S + 7)
(J + 7)
:
In 7 years, Scott will be twice as old as Sun Jung.
(S+7) = 2(J+7)
S + 7 = 2J + 14
S = 2J + 14 - 7
S = 2J + 7
:
Give the current age of Scott and Sun Jung.
Substitute (J+22) for S in the above equation and find J
J + 22 = 2J + 7
22 - 7 = 2J - J
J = 15 yrs now
:
Given that S is 22 yrs older:
S = 15 + 22 = 37 yrs age now
:
:
Check our solutions in the statement:
"In 7 years, Scott will be twice as old as Sun Jung."
37 + 7 = 2(15 + 7)
44 = 2(22); confirms our solution